# A Physics-Inspired Optimizer: Velocity Regularized Adam

### v1

@@ Abstract
We introduce Velocity-Regularized Adam (VRAdam), a physics-inspired optimizer for training deep neural networks that draws on ideas from quartic terms for kinetic energy with its stabilizing effects on various system dynamics. Previous algorithms, including the ubiquitous Adam, operate at the so called adaptive edge of stability regime during training leading to rapid oscillations and slowed convergence of loss. However, VRAdam adds a higher order penalty on the learning rate based on the velocity such that the algorithm automatically slows down whenever weight updates become large. In practice, we observe that the effective dynamic learning rate shrinks in high-velocity regimes, damping oscillations and allowing for a more aggressive base step size when necessary without divergence. By combining this velocity-based regularizer for global damping with Adam’s per-parameter scaling to create a hybrid optimizer, we demonstrate that VRAdam consistently exceeds the performance against standard optimizers including AdamW. We benchmark various tasks such as image classification, language modeling, image generation and generative modeling using diverse architectures and training methodologies including Convolutional Neural Networks (CNNs), Transformers, and GFlowNets.
@@ 1 Introduction
Optimizing the parameters of deep neural networks remains a cornerstone of progress in machine learning. Improving on the core idea of Stochastic Gradient Descent (SGD) Sutskever et al. (2013), adaptive methods like Adam (Adaptive Moment Estimation) Kingma and Ba (2017) have become ubiquitous due to their practical effectiveness across diverse tasks and architectures. Despite its success, the performance of Adam can be sensitive to hyperparameter choices and its training dynamics can exhibit instabilities Reddi et al. (2019). Furthermore, fully understanding the training dynamics of deep neural networks remains an open challenge Wang and Choromanska (2025), and even small improvements to existing optimization algorithms can often lead to significant reductions in resource consumption.
One line of work, observed empirically, is that training often occurs at the edge of stability Cohen et al. (2022, 2024), a regime for which the largest eigenvalue, also called sharpness, of the loss Hessian equilibrates around a fixed value proportional to the inverse of the learning rate (LR). This seems in contrast with common presumptions in classical optimization theory and has profound implications for convergence speed, stability, and generalization. In classical optimization, higher LRs lead to faster convergence at the cost of oscillations or divergence if stability constraints (depending on the loss landscape) are violated Boyd and Vandenberghe (2004). This pathological behaviour of optimizers, like AdamW, leads to instabilities and slowed convergence of the loss.
@@
This work introduces a class of physics-inspired optimizers, termed Velocity-Regularized Adam (VRAdam), designed to improve upon the stability and performance of standard adaptive methods Kingma and Ba (2017); Loshchilov and Hutter . Inspired by quartic terms used to model kinetic energy in more stable systems such as classical time crystals Shapere and Wilczek (2012) and heavy quark modeling using non-relativistic quantum chromodynamics (NRQCD) Braaten (1997) known for their unique stability properties, VRAdam introduces a novel regularization mechanism. This mechanism controls the effective learning rate $\eta$ via penalizing high velocity, namely $\eta_{t}=\alpha_{0}/(1+\min(\beta_{3}||v_{t}||^{2},\alpha_{1}))$. Crucially, this regularization can be configured to depend on either the norm of the optimizer’s velocity $||v_{t}||$ (i.e., the momentum buffer) or the norm of the gradient itself.
Equipped with this new optimizer, VRAdam, we probe its dynamics at the adaptive edge of stability, observe faster convergence, and analyze its sharpness. With well-tuned hyperparameters, we benchmark VRAdam against AdamW, SGD with Nesterov momentum and RMSProp on image classification with the CIFAR-10 dataset and convolutional neural network architecture, on language modeling with transformers on the WikiText2 dataset and a generative modeling task with GFlowNets and report improved performance on all tasks by up to 63%. Furthermore, we report low hyperparameter sensitivity on a diffusion model trained on the handwritten digit dataset, MNIST.
@@
  • Adaptive edge of stability analysis of VRAdam with faster convergence,
@@ 2 Background
This adaptive preconditioning coincides with the learning rate scaling in Adam (see Alg. 1) and scales down the step size in high-variance (typically high-curvature) directions as well as scales up in low-variance ones. Since the local stability of an optimizer around a minimizer depends on the eigenvalues of the quadratic Taylor approximation $L(x)\approx\frac{1}{2}x^{\top}Hx$, Adam’s dynamics are shown to be stable Cohen et al. (2024) as long as
@@
However, as this threshold is attained, the adaptive oscillatory regime can slow final convergence, as the optimizer continually adjusts its preconditioner to maintain stability as well Song and Yun (2023).
Physical origins of exotic Lagrangians. To better comprehend and navigate the edge of stability regime, we can draw inspiration from the physics governing complex optimization scenarios. The deep insights provided by the interplay of physics and machine learning frameworks have been demonstrated in various scenarios, such as the improved interpretation of Neural Tangent Kernels (NTK) through Langevin dynamics Avidan et al. (2025). One central concept in physics is the Lagrangian $\mathcal{L}(x,v)=T(v)-V(x)$ of a system, which is a function of position $x$ and velocity $v$ and (typically) defined as the difference between kinetic energy $T$ and potential energy $V$ from which the equation of motion can be derived via the Euler-Lagrange equation. In this work, we investigate non-standard Lagrangian formulations of physical phenomena with excellent stability conditions. For example, the stability of seemingly disparate quantum systems like heavy quarkonia (described by NRQCD) and classical time crystals share conceptual parallels rooted in higher-order velocity terms. These terms fundamentally reshape energy landscapes by creating non-standard dispersion relations, establishing invariant submanifolds in phase space where stable configurations emerge as attractors or limit cycles. By explicitly breaking certain symmetries—Lorentz invariance in NRQCD and time-translation in time crystals—these higher-order kinetic terms paradoxically enhance stability through topological protection mechanisms and the generation of emergent length/time scales Niemi (2021); Guha and Ghose-Choudhury (2019).
As a demonstration of this phenomenon, we can consider Nonrelativistic QCD (NRQCD), which is an effective field theory that expands full quantum chromodynamics in inverse powers of a heavy-quark mass $m$ Assi et al. (2023). The bilinear heavy-quark Lagrangian $\mathcal{L}$ up to $O(1/m^{3})$ reads,
$\mathcal{L}_{\mathrm{NRQCD}}=\psi^{\dagger}\left(iD_{0}-m+\frac{\mathrm{D}^{2}}{2m}+\frac{\mathrm{D}^{4}}{8m^{3}}+c_{F}\frac{g\sigma\mathrm{~{}B}}{2m}+\cdots\right)\psi,$ (3)
where $D$ represents the covariant derivative, $\psi$ is the heavy quark field, and $\frac{g\sigma\mathrm{~{}B}}{2m}$ is the term associated the chromomagnetic field interactions. Crucially, the $\frac{\mathrm{D}^{4}}{8m^{3}}$ operator acts as the leading positive quartic kinetic correction, ensuring the heavy-quark Hamiltonian remains bounded and stable with relativistic expansions Brambilla et al. (2000). Further details on the stability provided by the quartic term is provided in Appendix A.
@@ 3 Method
To translate this physics insight into an optimizer design, we identify the stabilizing aspects of such phenomena such as the heavy-quark momentum with the optimizer’s global momentum buffer $v$. Accordingly, we posit a kinetic energy of the form: $T_{\mathrm{VRAdam}}(v)=\frac{m}{2}\|v\|^{2}+\frac{\beta_{3}}{4}\|v\|^{4}$, where $m$ is the mass and $\beta_{3}$ is a tunable parameter. The Lagrangian then becomes:
@@
where we solve for Euler-Lagrange equation: $\frac{d}{dt}\frac{\partial\mathcal{L}}{\partial v}-\frac{\partial\mathcal{L}}{\partial x}=0$. Such that,
$\frac{\partial\mathcal{L}}{\partial x}=-\frac{\partial V(x)}{\partial x}=-\nabla L_{\text{loss }}(x).$ (5)
We know that the loss of a neural network with parameters $x$ can be thought of as the potential landscape Holderrieth et al. (2024).
$\begin{gathered}\frac{\partial\mathcal{L}}{\partial v}=\frac{\partial T_{\mathrm{VRAdam}}}{\partial v}=\frac{\partial}{\partial v}\left(\frac{1}{2}m\|v\|^{2}+\frac{1}{4}\beta_{3}(\|v\|^{2})^{2}\right)\\ \frac{\partial\mathcal{L}}{\partial v}=\left(m+\beta\|v\|^{2}\right)v\end{gathered}$ (6)
The generalized momentum in this form is given by:
$p=\frac{\partial L}{\partial v}=\left(m+\beta_{3}\|v\|^{2}\right)v$ (7)
The time derivative for the generalized momentum is then
$\frac{d}{dt}\frac{\partial L}{\partial v}=\frac{d}{dt}\left[\left(m+\beta_{3}\|v\|^{2}\right)v\right]$ (8)
@@
$\dot{v}=-\nabla L_{\text{Loss}}(x)/(m+3\beta_{3}||v||^{2}),\dot{x}=v$ (10)
where the dot corresponds to the time derivative. Rather than explicitly constructing an optimizer based on an ordinary differential equation solver for Eq. 10 (via discretization and introduction of dissipation, e.g. França et al. (2020)), we utilize the term, $(1/(m+3\beta_{3}||v||^{2})$, and embed it as a dynamic learning rate into AdamW, to enhance the successfully proven properties. In particular, this approach avoids choosing specific forms associated with various different integrators. The full velocity regularized Adam is given in Alg. 1, where we highlight in blue the changes to AdamW.
@@
For VRAdam, we obtain, through reparameterization and modification, the dynamic learning rate $\eta_{t}=\alpha_{0}/(1+\min(\beta_{3}||v_{t}||^{2},\alpha_{1}))$ for timestep $t$, where $\alpha_{0}$ and $\alpha_{1}$ control the maximal and minimal LR respectively,and $\beta_{3}$ as the velocity penalizer. This is inspired by the bound introduced to $v^{2}$ in physical setting as discussed in Appendix A. The parameterization of LR, compared to the physically derived one, clips the velocity to avoid getting stuck if gradients and therefore velocity become large. Furthermore, replacing the norm of the velocity (exponential moving average of gradients) with the current gradient norm (a more local property) can lead to improved performance in some scenarios. Weight decay is applied in the traditional manner.
@@ 4 Analysis
Here, we analyze the behavior of VRAdam in the adaptive edge of stability regime compared to that of Adam. For this analysis, following Cohen et al. (2024), we train a ResNet 32 architecture on CIFAR-10 for an image classification task, both with VRAdam and Adam. The training is stopped as soon as the training loss falls below 0.1 or an accuracy of 0.97 is reached. In Fig. 2 (a) and (b), the training curves of VRAdam indicate faster convergence in both minimizing training loss as well as maximizing training accuracy as compared to Adam. The training curves are smooth for both optimizers. When juxtaposed with the sharpness comparison depicted in Fig. 2 (c), we observe that the maximum eigenvalue (sharpness) of the preconditioned Hessian of the loss, remains consistently lower due to the dynamic learning rate adjustments induced by VRAdam. We can empirically observe that this “self-braking mechanism ” enables VRAdam to operate in a regime of optimal lower sharpness and escape sharp oscillations, leading to faster convergence. (The negative eigenvalues are likely artifacts of the power method used to compute the largest eigenvalue (Trefethen and Bau, 1997).)
@@
The effective learning rate of VRAdam is shown in Fig 2 (d). During the first 25 iterations, the learning rate dynamically increases close to the maximal value allowed LR and then proceeds to decrease while exhibiting oscillatory behavior as we converge to the minima. The bound on the minimal LR is not active in this example, while the base LR of Adam stays constant throughout training. As described in the seminal work of Schmidhuber et al., we note that minima with lower sharpness are associated with better generalization Hochreiter and Schmidhuber (1997); Foret et al. (2021). We can also observe, that VRAdam’s dynamic learning rate quickly moves to the maximal LR to exploit the loss landscape optimally, while exploiting the trade-off between the adaptive edge of stability and faster convergence.
@@
Figure 2: (a) Training loss curves for VRAdam and Adam of ResNet 32 on CIFAR-10 (b) training accuracy curves (c) plot of maximal eigenvalues of the loss Hessian d effective learning rate during training. Hyperparameters for these plots are provided in Appendix B.5.

### v2

@@ Abstract
We introduce Velocity-Regularized Adam (VRAdam), a physics-inspired optimizer for training deep neural networks that draws on ideas from quartic terms for kinetic energy with its stabilizing effects on various system dynamics. Previous algorithms, including the ubiquitous Adam, operate at the so-called adaptive edge of stability regime during training, leading to rapid oscillations and slowed convergence of loss. However, VRAdam adds a higher order penalty on the learning rate based on the velocity such that the algorithm automatically slows down whenever weight updates become large. In practice, we observe that the effective dynamic learning rate shrinks in high-velocity regimes, and damping oscillations. By combining this velocity‑based regularizer for global damping with Adam’s per‑parameter scaling, we create a powerful hybrid optimizer. For this optimizer, we provide rigorous theoretical analysis of operation at the edge of stability from a physical and control perspective for the momentum. Furthermore, we derive convergence bounds with the rate $\mathcal{O}(\ln(N)/\sqrt{N})$ for a stochastic non‑convex objective under mild assumptions. We demonstrate that VRAdam exceeds the performance against standard optimizers including AdamW. We benchmark various tasks such as image classification, language modeling, and generative modeling using diverse architectures and training methodologies including Convolutional Neural Networks (CNNs), Transformers, and GFlowNets.
@@ 1 Introduction
Optimizing the parameters of deep neural networks remains a cornerstone of progress in machine learning. Improving on the core idea of Stochastic Gradient Descent (SGD) Sutskever et al. (2013), adaptive methods like Adam (Adaptive Moment Estimation) Kingma & Ba (2017) have become ubiquitous due to their practical effectiveness across diverse tasks and architectures. Despite its success, the performance of Adam can be sensitive to hyperparameter choices and its training dynamics can exhibit instabilities Reddi et al. (2019). Furthermore, fully understanding the training dynamics of deep neural networks remains an open challenge Wang & Choromanska (2025), and even small improvements to existing optimization algorithms can often lead to significant reductions in resource consumption.
One line of work, observed empirically, is that training often occurs at the edge of stability Cohen et al. (2022; 2024), a regime for which the largest eigenvalue, also called sharpness, of the loss Hessian equilibrates around a fixed value proportional to the inverse of the learning rate (LR). This seems in contrast with common presumptions in classical optimization theory and has profound implications for convergence speed, stability, and generalization. In classical optimization, higher LRs lead to faster convergence at the cost of oscillations or divergence if stability constraints (depending on the loss landscape) are violated Boyd & Vandenberghe (2004). This pathological behavior of optimizers, like AdamW, leads to instabilities and a slowed decrease of the loss.
@@
[…]
Equipped with this new optimizer, VRAdam, we probe its dynamics at the adaptive edge of stability, observe faster convergence, and analyze its sharpness empirically against AdamW and Sharpness Aware Minimization (SAM) Foret et al. (2021). We also introduce rigorous theoretical analysis of the, global uniform exponential stability of momentum, Weber et al. (2024), as well as a physical inspired Lyapunov candidate derived from our Lagrangian that demonstrates stability properties. With well-tuned hyperparameters, we benchmark VRAdam against AdamW, RAdam Liu et al. (2020), SGD with Nesterov momentum and RMSProp Ruder (2017) on image classification with the CIFAR-10 dataset and the convolutional neural network architecture, on language modeling with transformers on the WikiText2 dataset, and a generative modeling task with GFlowNets and report improved performance on all tasks. We also report increased target-to-loss against AdamW on large scale training for language models such as GPT Brown et al. (2020) with marginal computational increase in overhead.
@@
  • Adaptive edge of stability analysis of VRAdam with faster convergence and associated empirical and theoretical evidence from momentum analysis,
  • Convergence bound for non-convex stochastic objective,
@@ 2 Background
This adaptive preconditioning coincides with the learning rate scaling in Adam (see Alg. 1) and scales down the step size in high‐variance (typically high‐curvature) directions as well as scales up in low‐variance ones. Since the local stability of an optimizer around a minimizer depends on the eigenvalues of the quadratic Taylor approximation $L(x)\approx\frac{1}{2}x^{\top}Hx$, Adam’s dynamics are shown to be stable Cohen et al. (2024) as long as
@@
However, as this threshold is attained, the adaptive oscillatory regime can slow final convergence, as the optimizer continually adjusts its preconditioner to maintain stability as well Song & Yun (2023).
Physical origins of exotic Lagrangians. To better comprehend and navigate the edge of stability regime, we can draw inspiration from the physics governing complex optimization scenarios. The deep insights provided by the interplay of physics and machine learning frameworks have been demonstrated in various scenarios, such as the improved interpretation of Neural Tangent Kernels (NTK) through Langevin dynamics Avidan et al. (2025). One central concept in physics is the Lagrangian $\mathcal{L}(x,v)=T(v)-V(x)$ of a system, which is a function of position $x$ and velocity $v$ and (typically) defined as the difference between kinetic energy $T$ and potential energy $V$ from which the equation of motion can be derived via the Euler-Lagrange equation. In this work, we investigate non-standard Lagrangian formulations of physical phenomena with excellent stability conditions. For example, the stability of seemingly disparate quantum systems like heavy quarkonia (described by NRQCD) and classical time crystals share conceptual parallels rooted in higher-order velocity terms. These terms fundamentally reshape energy landscapes by creating non-standard dispersion relations, establishing invariant submanifolds in phase space where stable configurations emerge as attractors or limit cycles. Further information regarding these systems are found in Appendix A.
@@ 3 Method
To translate this physics insight into an optimizer design, we identify the stabilizing aspects of such phenomena such as the heavy‐quark momentum with the optimizer’s global momentum buffer $v$. Accordingly, we posit a kinetic energy of the form: $T_{\mathrm{VRAdam}}(v)=\frac{m}{2}\|v\|^{2}+\frac{\beta_{3}}{4}\|v\|^{4}$, where $m$ is the mass and $\beta_{3}$ is a tunable parameter. The Lagrangian then becomes
@@
for which we solve the Euler-Lagrange equation $\frac{d}{dt}\frac{\partial\mathcal{L}}{\partial v}-\frac{\partial\mathcal{L}}{\partial x}=0$. We know that the loss of a neural network with parameters $x$ can be thought of as the potential landscape Holderrieth et al. (2024), such that $\frac{\partial\mathcal{L}}{\partial x}=-\frac{\partial V(x)}{\partial x}=-\nabla L_{\text{loss }}(x)$.
@@
[…]
where the dot corresponds to the time derivative. Rather than explicitly constructing an optimizer based on an ordinary differential equation solver for Eq. 5 (via discretization and introduction of dissipation, e.g. França et al. (2020)), we utilize the term, $(1/(m+3\beta_{3}||v||^{2})$, and embed it as a dynamic learning rate into AdamW, to enhance the successfully proven properties. Note, that this expression from Eq. 5 is derived for the 1-D case, where the dynamics $\dot{v}$ is collinear with $v$. In particular, this approach avoids choosing specific forms associated with various different integrators. The full velocity regularized Adam is given in Alg. 1, where we highlight in blue the changes to AdamW.
@@
[…]
@@ 4.1 Empirical Analysis
For this analysis, following Cohen et al. (2024), we train a ResNet 32 architecture on CIFAR-10 for an image classification task, with VRAdam, Adam and SAM. The training is stopped as soon as the training loss falls below 0.1 or an accuracy of 0.97 is reached. In Fig. 2 (a) and (b), the training curves of VRAdam indicate faster convergence in both minimizing training loss as well as maximizing training accuracy as compared to Adam and SAM, known for sharpness minimization. The training curves are smooth for both optimizers. When juxtaposed with the sharpness comparison depicted in Fig. 2 (c), we observe that the maximum eigenvalue (sharpness) of the preconditioned Hessian of the loss, remains adaptable to faster convergence due to the dynamic learning rate adjustments induced by VRAdam.
@@
The effective learning rate of VRAdam is shown in Fig. 2 (d). During the first 25 iterations, the learning rate dynamically increases close to the maximal value allowed LR and then proceeds to decrease while exhibiting oscillatory behavior as we converge to the minima. The bound on the minimal LR is not active in this example, while the base LR of Adam stays constant throughout training. As described in the seminal work of Schmidhuber et al., we note that minima with lower sharpness are associated with better generalization Hochreiter & Schmidhuber (1997); Foret et al. (2021). We can also observe, that VRAdam’s dynamic learning rate quickly moves to the maximal LR to exploit the loss landscape optimally, while exploiting the trade-off between the adaptive edge of stability and faster convergence.
@@
Figure 2: (a) Training loss curves for VRAdam, Adam, and SAM Foret et al. (2021) of ResNet 32 on CIFAR-10 (b) training accuracy curves (c) plot of maximal eigenvalues of the loss Hessian d effective learning rate during training. Hyperparameters for these plots are provided in Appendix D.4.
4.2 Stability of VRAdam
We analyze the behavior of VRAdam in the adaptive edge of stability regime compared to that of Adam in a momentum ablation setting without parameter-scaling based on second order moments estimates, bias corrections or weight decay. By simply adding weight decay, the admissible base step is constrained, the rest of the proof remains as follows.
Objective Function: Traditionally for stability analysis, we consider a convex quadratic objective:
$f(\theta)=\frac{1}{2}(\theta-\theta^{*})^{\top}H(\theta-\theta^{*})$
where $H$ is a positive definite Hessian matrix. Let $x_{t}=\theta_{t}-\theta^{*}$ be the error. The gradient is, $g_{t}=\nabla f(\theta_{t-1})=Hx_{t-1}$. Let $\mu=\lambda_{\min}(H)$ and $L:=\lambda_{\max}(H)$. From Algorithm 1, we know that $\eta_{t}\in\mathcal{H}:=\left[\eta_{\min},\alpha_{0}\right]$ and $\eta_{\min}=\alpha_{0}/\left(1+\alpha_{1}\right)>0$. Let $\beta_{1}$ from Algorithm 1 be $\beta$ for this analysis. The following theorem and proof provides sufficient but not necessary conditions for stability in discrete time. Thus proving a more restrictive case of the optimizer.
Theorem 4.1 (Uniform exponential stability of VRMomentum).
Consider $f(\theta)$ with $0\prec H\preceq LI$. Let VRMomentum be $\beta\in[0,1)$, $\beta_{3}>0$, $\alpha_{0}>0$, $\alpha_{1}\in(0,\infty]$, no bias correction, and $\lambda\geq 0$. If $\alpha_{0}L<B(\beta)=\frac{2(1+\beta)}{1-\beta}$ or $\eta_{\min}L<B(\beta)$ if the LR clip is active, then for any realization of $\{\eta_{t}\}$ generated by the gate $\eta_{t}=\alpha_{0}/(1+\min(\beta_{3}\|v_{t}\|^{2},\alpha_{1}))$, the origin is a globally uniformly exponentially stable equilibrium. Moreover, there exists a Common Quadratic Lyapunov Function (CQLF), $V(z)=z^{\top}Pz$ with $P\succ 0$ such that $V(z_{t})\leq(1-\epsilon)V(z_{t-1})$ for some $\epsilon\in(0,1)$.
Proof. Define the 2d-state $z_{t}=(x_{t},v_{t})$ and the parameterized update matrix
$A(\eta):=\begin{pmatrix}I-\eta(1-\beta)H&-\eta\beta I\\ (1-\beta)H&\beta I\end{pmatrix},\quad z_{t}=A(\eta_{t})z_{t-1}.$ (6)
Let $H=Q^{\top}\text{diag}(h_{i})Q$ with $0<\mu\leq h_{i}\leq L$. In the eigenbasis of $H$, the dynamics split into $d$ identical $2\times 2$ subsystems that share the same scalar $\eta_{t}$ and differ only by the curvature $h\in[\mu,L]$:
$\begin{pmatrix}\xi_{t}\\ v_{t}\end{pmatrix}=A_{h}(\eta_{t})\begin{pmatrix}\xi_{t-1}\\ v_{t-1}\end{pmatrix},\quad A_{h}(\eta):=\begin{pmatrix}1-\eta(1-\beta)h&-\eta\beta\\ (1-\beta)h&\beta\end{pmatrix}$. (7)
Hence it suffices to build a CQLF for the family $\{A_{h}(\eta):h\in[\mu,L],\eta\in\mathcal{H}\}$.
For a fixed $(h,\eta)$, the characteristic polynomial is
$\lambda^{2}-(1+\beta-\eta(1-\beta)h)\lambda+\beta=0,$ (8)
so the Schur criterion gives stability iff $\eta(1-\beta)h<2(1+\beta)$. In particular, if
$\alpha_{0}L<B:=\frac{2(1+\beta)}{1-\beta}$ (A1)
then every $A_{h}(\eta)$ with $\eta\in\mathcal{H}$ is Schur-stable. (This is the Adam/AEoS bound specialized to $P=I$ Cohen et al. (2024))
We now produce $P\succ 0$ such that
$A_{h}(\eta)^{\top}PA_{h}(\eta)-P\preceq-\epsilon I_{2}\quad\forall\eta\in\mathcal{H},\forall h\in[\mu,L],$ (A2)
for some $\epsilon>0$. Take the block-diagonal, curvature-agnostic form
$P=\text{diag}(p_{1}I,p_{2}I),\quad p_{1}>0,p_{2}>0,$ (A3)
which lifts to $P_{d}=\text{diag}(p_{1}I_{d},p_{2}I_{d})$ in $2d$ dimensions.
For a fixed $(h,\eta)$, set
$\Delta(\eta,h):=A_{h}(\eta)^{\top}PA_{h}(\eta)-P=\begin{pmatrix}\Delta_{11}&\Delta_{12}\\ \Delta_{12}&\Delta_{22}\end{pmatrix}$ (9)
with
$\Delta_{11}(\eta,h)$ $=p_{1}(-2\eta(1-\beta)h+\eta^{2}(1-\beta)^{2}h^{2})+p_{2}(1-\beta)^{2}h^{2},$ (10)
$\Delta_{22}(\eta,h)$ $=p_{1}\eta^{2}\beta^{2}+p_{2}(\beta^{2}-1),$ (11)
$\Delta_{12}(\eta,h)$ $=-p_{1}\eta\beta(1-\eta(1-\beta)h)+p_{2}\beta(1-\beta)h.$ (12)
We first force strict negativity on the diagonal, uniformly over $\eta\in\mathcal{H}$ and $h\in[\mu,L]$. Using the bounds $\eta\in[\eta_{\min},\alpha_{0}]$ and $h\in[\mu,L]$, a sufficient pair of conditions for $\Delta_{11}\leq-\delta_{1}$ and $\Delta_{22}\leq-\delta_{2}$ is
$\underbrace{(\alpha_{0}^{2}p_{1}+p_{2})L}_{\text{upper bound for the $h^{2}$-term}}$ $<\underbrace{\frac{2\eta_{\min}}{1-\beta}p_{1}}_{\text{lower bound for the linear $h$-term}},$ (A4)
$\underbrace{p_{2}(1-\beta^{2})}_{\text{negative term magnitude}}$ $>\underbrace{p_{1}\alpha_{0}^{2}\beta^{2}}_{\text{positive term bound}}.$ (A5)
(Here we used the worst cases $\eta=\alpha_{0}$ and $h=L$ wherever they make the expression largest.)
The pair (A4)–(A5) admits a solution $(p_{1},p_{2})$ iff
$\alpha_{0}^{2}L<\frac{2(1+\beta)}{\eta_{\min}}\quad\iff\quad\eta_{\min}>\frac{\alpha_{0}^{2}L}{2(1+\beta)}$ (A6)
Indeed, set $p_{1}=1$ and choose any
$p_{2}\in\left(\frac{\alpha_{0}^{2}\beta^{2}}{1-\beta^{2}},\frac{2\eta_{\min}}{(1-\beta)L}-\alpha_{0}^{2}\right).$ (A7)
The interval is nonempty precisely when (A6) holds.
Next, to pass from diagonal negativity to matrix negativity, we use the Schur complement:
$\Delta\preceq-\epsilon I_{2}\iff\Delta_{11}\leq-\epsilon,\quad\Delta_{22}-\frac{\Delta_{12}^{2}}{\Delta_{11}}\leq-\epsilon.$ (13)
Under (A4)–(A5), $\Delta_{11}$ and $\Delta_{22}$ are uniformly $\leq-\delta$ for some $\delta>0$. A direct (but routine) bound gives, for all $\eta\in\mathcal{H},h\in[\mu,L]$,
$\frac{\Delta_{12}^{2}}{-\Delta_{11}}\leq\frac{\beta^{2}(p_{1}\alpha_{0}+p_{2}(1-\beta)L)^{2}}{2\eta_{\min}p_{1}-(\alpha_{0}^{2}p_{1}+p_{2})(1-\beta)L},$ (14)
and the right-hand side is strictly smaller than $\Delta_{22}$ when (A4)–(A5) hold with slack. Thus there exists $\epsilon>0$ such that (A2) is satisfied.
For fixed $h$ and $P\succ 0$, the map
$\eta\mapsto A_{h}(\eta)^{\top}PA_{h}(\eta)=(A_{0}+\eta A_{1})^{\top}P(A_{0}+\eta A_{1})$ (15)
is matrix-convex in $\eta$ (its second derivative is $2A_{1}^{\top}PA_{1}\succeq 0$). Therefore, if the inequality
$A_{h}(\eta)^{\top}PA_{h}(\eta)-P\preceq-\epsilon I_{2}$ (16)
holds at the endpoints $\eta=\eta_{\min}$ and $\eta=\alpha_{0}$, it holds for all $\eta\in[\eta_{\min},\alpha_{0}]$. Under (A6), the choice (A7) does exactly that.
The block choice $P_{d}=\text{diag}(p_{1}I_{d},p_{2}I_{d})$ certifies
$A(\eta)^{\top}P_{d}A(\eta)-P_{d}\preceq-\epsilon I_{2d}\quad\forall\eta\in\mathcal{H},$ (17)
hence the quadratic Lyapunov $V(z)=z^{\top}P_{d}z$ yields
$V(z_{t})\leq(1-c)V(z_{t-1}),\quad c=\frac{\epsilon}{\lambda_{\max}(P_{d})}\in(0,1),$ (18)
which implies global uniform exponential stability and concludes the proof.
In the Appendix B, an alternative Lyapunov candidate based on Lagrangian physics is presented.
4.3 Convergence Analysis
In this section, we provide a convergence analysis of VRAdam by extending the work of Défossez et al. (2022) on Adam. The main result is a probabilistic bound that VRAdam converges in expectation to a stationary point for a nonconvex (and therefore also a convex) objective function under a set of regularity assumptions and simplifications.
Setup: The objective function is given by $F(\theta):\mathbb{R}^{d}\to\mathbb{R}$. $\nabla f$ are gradient estimates of $F$ such that $\mathbb{E}[\nabla f]=\nabla F$. The random variable $\tau_{N}$ describes an index with probabilities given by $\mathbb{P}[\tau=j]\propto 1-\beta_{1}^{N-j}$ for $j\in\{0,1,...,N-1\}$. The constants $\alpha_{\min}=\alpha_{0}/(1+\alpha_{1})>0$ and $\alpha_{\max}=\alpha_{0}>0$ correspond to the minimal and maximal possible learning rate at every step. $F$ is bounded from below, namely $\forall\theta,F(\theta)\geq F_{*}$. The gradient estimates are almost surely bounded in the $\ell_{\infty}$-norm, namely $\left\|\nabla f\right\|_{\infty}\leq R-\sqrt{\epsilon}$. $\nabla F$ is L-Lipschitz-continuous in the $\ell_{2}$-norm, namely $\forall\theta,\phi,\left\|\nabla F(\theta)-\nabla F(\phi)\right\|_{2}\leq L\left\|\theta-\phi\right\|_{2}$. The weight decay parameter $\lambda=0$ is set to zero. The bias correction for the velocity $\hat{v}_{t}\leftarrow v_{t}/(1-\beta_{1}^{t})$ is not included. The hyperparameter $\beta_{1},\beta_{2}$ fulfill $0<\beta_{2}<1$, $0\leq\beta_{1}<\beta_{2}$ and the number of iterates satisfies $N>\beta_{1}/(1-\beta_{1})$.
Theorem 4.2 (Convergence of VRAdam).
Given the six assumptions above, the iterates $\theta_{t}$ of Algorithm 1 fulfill:
$\mathbb{E}\bigl[\|\nabla F(\theta_{\tau})\|^{2}\bigr]\leq 2R\,\frac{F(\theta_{0})-F_{*}}{\alpha_{\min}\,\tilde{N}}\\ \quad+\,E\Bigl(\tfrac{1}{\tilde{N}}\ln\bigl(1+\tfrac{R^{2}}{(1-\beta_{2})\epsilon}\bigr)-\tfrac{N}{\tilde{N}}\ln(\beta_{2})\Bigr).$
with $\tilde{N}=N-\frac{\beta_{1}}{1-\beta_{1}},\quad\text{and}$
$E=\;\frac{\alpha_{\max}^{2}dRL(1-\beta_{1})}{\alpha_{\min}(1-\beta_{1}/\beta_{2})(1-\beta_{2})}+\;\frac{12\,\alpha_{\max}dR^{2}\sqrt{1-\beta_{1}}}{\alpha_{\min}(1-\beta_{1}/\beta_{2})^{3/2}\,\sqrt{1-\beta_{2}}}+\;\frac{2\,\alpha_{\max}^{3}dL^{2}\beta_{1}}{\alpha_{\min}(1-\beta_{1}/\beta_{2})\,(1-\beta_{2})^{3/2}}.$
A proof can be found in the appendix C. The right-hand side of the bound converges to zero in the limit of $N\to\infty$ if $\alpha_{min}(N)\propto N^{-1/2}$, $\alpha_{max}(N)\propto N^{-1/2}$, and $1-\beta_{2}(N)\propto N^{-1}$ to leading order. For this scaling of hyperparameters, VRAdam has the same convergence rate of $\mathcal{O}(\ln(N)/\sqrt{N})$ as Adam. While this convergence result provides a strong theoretical foundation, benchmarking is key to understanding performance under realistic conditions.